Circular No. 6465

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) L. H. Wasserman, Lowell Observatory, writes that comet C/1995 O1

will occult the star PPM 200723 (α = 17^{h}23^{m}59^{s}.845, δ = −4^{o}48′09′′.45, equinox 2000.0, epoch 1996.8; V = 9.2) on Oct. 5.13 UT, the nominal track currently crossing the U.S. from southern California to northeastern Minnesota. The comet has a sky-plane velocity of 0′′.092 (5.1 km)/s. Detailed predictions and a map of the ground track are available at http://www.

lowell.edu/users/buie/occ/predict.html.

Total visual magnitudes (moonlight interfering during Aug. 23–27): Aug. 17.25 UT, 5.1 (R. Keen, Mt. Thorodin, CO, naked eye); 19.10, 5.7 (J. Bortle, Stormville, NY, 10 × 50 binoculars; tail 1^{o}.0); 21.22, 5.4 (C. E. Spratt, Victoria, BC, 7 × 50 binoculars); 23.87, 5.6 (A. Pereira, Carnaxide, Portugal, 9 × 34 binoculars); 25.84, 5.0 (K. Hornoch, Lelekovice, Czech Republic, 10 × 50 binoculars); 27.87, 5.6 (Pereira); 29.87, 5.7 (Pereira).

Corrigendum. On IAUC 6463, line 15, for F. Crippa read R. Crippa

COMET C/1996 P2 (RUSSELL-WATSON)

Improved parabolic orbital elements from MPC 27690:

T = 1996 Feb. 29.518 TT  ω = 188.^{o}661
Ω = 101.556 }2000.0
q = 2.01067 AU  i = 29.134
1996TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Sept. 4 1^{h}38.^{m}58 −33^{o}42.^{′}8 2.183 2.969 133.^{o}4 14.^{o}3 12.4
9 1 36.36 −34 21.3 2.206 3.009 135.5 13.6 12.5
14 1 33.63 −34 54.8 2.234 3.049 137.2 13.0 12.6
19 1 30.49 −35 22.3 2.266 3.090 138.4 12.5 12.7
24 1 27.03 −35 43.4 2.304 3.131 139.0 12.1 12.8
29 1 23.38 −35 57.6 2.347 3.172 139.1 11.9 12.9
Oct. 4 1 19.62 −36 04.7 2.395 3.214 138.5 11.9 13.0
9 1 15.87 −36 04.7 2.449 3.255 137.4 12.0 13.1
14 1 12.24 −35 57.6 2.508 3.297 135.8 12.2 13.2

Total visual magnitude estimates: Aug. 17.13, 12.2 (J. M. San Juan, Teruel, Spain, 0.44-m reflector); 19.08, 12.8 (J. Carvajal, Teruel, Spain, 0.44-m reflector); 21.43, 12.7 (A. Hale, Cloudcroft, NM, 0.41-m reflector); 25.02, 12.3 (K. Sárneczky, Ráktanya, Hungary, 0.44-m reflector).

1996 August 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Brian G. Marsden